OpenStudy (anonymous):

What is the surface area of a sphere if the volume is 1928pie

OpenStudy (mathstudent55):

\(V = \dfrac{4}{3}\pi r^3\) Let V equal the given volume, and solve for r. Once you find r, the formula for the area of a sphere is: \(A = 4 \pi r^2\)